# Mokav: Execution-driven Differential Testing with LLMs

### v1

@@ 6. Threats to Validity
Construct Validity: The main threat to the construct validity of this study is data leakage in LLMs (Al-Kaswan et al., 2024). As LLMs are trained on a large corpus of open-source data, they may have some program pairs of our datasets in their training data. This may threaten the generalizability of our experimental conclusions. Despite this concern, we note open-source repositories usually do not put difference-exposing tests next to programs. This means there is no direct connection between programs and DETs in LLMs’ training data. Consequently, the data leakage problem is less significant in DET generation, the goal of Mokav, compared to more traditional tasks, such as code generation (Du et al., 2024) and program repair (Jiang et al., 2023). Also, to further alleviate the data leakage concern, we assess the effectiveness of Mokav on two different datasets, i.e., QuixBugs (Lin et al., 2017) and C4DET.
@@
Internal Validity: Mokav uses a high temperature for the LLM in its default configuration. This leads to randomness in its responses, a threat to the internal validity of our experimental results. To address this threat, we repeat our main experiment of generating DETs for program pairs in C4DET five times. The results show that in all five instances, the effectiveness of Mokav is between 80.1% and 81.7%. This indicates that the main conclusions of our study are robust and only slightly affected by the randomness of LLMs.
7. Related Work
@@
Test Generation with LLMs LLMs are being applied on many software engineering tasks (Fan et al., 2023), and test generation is one of the major targets (Wang et al., 2024; Siddiq et al., 2023; Xie et al., 2023; Tang et al., 2024; Lemieux et al., 2023; Tufano et al., 2020; Altmayer Pizzorno and Berger, 2024; Chen et al., 2022; Takerngsaksiri et al., [n. d.]; Plein et al., 2023; Lahiri et al., 2022; Vikram et al., 2023; Liu et al., 2024a; Nie et al., 2023; Ryan et al., 2024; Dakhel et al., 2024; Pizzorno and Berger, 2024; Alshahwan et al., 2024; Steenhoek et al., 2023). A test consists of multiple parts: an input to a program under test (PUT), a test setup that moves the PUT into a specific state, and an oracle that asserts the PUT correctly produces the expected outputs. LLMs are used to generate all these parts of tests.
@@
Deng et al. (Deng et al., 2023) introduce TitanFuzz. This tool employs LLMs to generate human-like code that tests the API of deep learning (DL) libraries. TitanFuzz uses LLMs to generate and mutate input DL programs for fuzzing. Xia et al. (Xia et al., 2024) propose Fuzz4All, an LLM-based fuzzer that generates input test programs for compilers. Fuzz4All iteratively prompts the LLM to create new test inputs. The authors show that Fuzz4All effectively generates test inputs for six different programming languages. While TitanFuzz and Fuzz4All generate input programs for a single system, Mokav generates test inputs that differentiate two given programs.
@@
LLMs have also been used for oracle generation (Watson et al., 2020; Tufano et al., 2022; Hossain and Dwyer, 2024; Li and Yuan, 2024). Dinella et al. propose TOGA (Dinella et al., 2022). TOGA uses a fine-tuned version of CodeBERT (Feng et al., 2020) to generate assertion candidates based on the PUT and test prefix and rank these candidate assertions. Liu et al. (Liu et al., 2023b) show limitations of TOGA: its unrealistic assumption of the availability of a correct focal method, its ignorance of important evaluation metrics, such as precision, and the lack of a straightforward baseline. Zhang et al. (Zhang et al., 2024) generate tests in which the oracle expects an exception. In contrast to these works, Mokav focuses on input generation instead of oracle generation.
@@
Li et al. (Li et al., 2023) introduce differential prompting, an LLM-based tool that generates fault-inducing tests. This tool takes a buggy program. It first uses the LLM to generate an intention description for the program. Then, it generates a fixed version of the program. Next, it employs the LLM to generate a list of test inputs. Finally, a test input that leads to different results on the buggy and fixed versions is labeled as fault-inducing. AID (Liu et al., 2024b) is another tool that follows the same procedure, while it uses the LLM to generate a test generator, instead of test inputs. These two tools are different from Mokav as they generate fault-inducing tests for a given buggy program. In contrast, Mokav directly generates difference exposing tests for two given programs.
@@
Liu et al. (Liu et al., 2024c) use LLMs to amplify existing tests in the HumanEval dataset (Chen et al., 2021). They generate new tests that expose differences between the dataset programs and incorrect programs generated for the code documentation. Their test generation method starts from test seeds generated by LLMs and then improves them by type-aware mutation. This method is different from Mokav, which uses an execution-driven iterative approach to improve initial tests generated by the LLM.
@@
Generating Tests for Python Generating tests for dynamically typed programming languages is challenging (Selakovic et al., 2018). Auger (Laffra, 2024) is one of the first tools to address this task for Python. Auger uses Python tracer to collect the values with which target functions are called and the values that target functions return. The collected data is then used to generate unit tests.
@@
The state-of-the-art tool for Python test generation is Pynguin (Lukasczyk and Fraser, 2022; Lukasczyk et al., 2020; Erni et al., 2024), which uses evolutionary algorithms to generate unit tests with high coverage. It uses type annotations and static analysis to infer the type of function parameters. Lukasczyk et al. (Lukasczyk et al., 2023) conduct a study to improve Pynguin’s initial type inference strategy and find a more suitable evolutionary algorithm: DynaMOSA (Panichella et al., 2017).
@@
Schoofs et al. (Schoofs et al., 2022) introduce AmPyfier, which amplifies existing Python test suites. AmPyfier runs existing tests on the PUT and collects values during execution. These collected values are then added to the existing test as new assertions. AmPyfier also amplifies the inputs to PUT iteratively. To make the amplification efficient, at each iteration, AmPyfier only keeps the assertion and input amplifications that increase mutation or code coverage score.
@@
Differential Testing Differential testing is a well-established technique for detecting anomalies and defects in programs (McKeeman, 1998; Evans and Savoia, 2007; Taneja and Xie, 2008b; Nilizadeh et al., 2019; Rutledge and Orso, 2022; Ye et al., 2023). It is employed to identify issues in many types of programs, such as robotic vehicle programs (Kim et al., 2023), game agents (Castellano et al., 2022), spreadsheet applications (Almasi et al., 2018), certificate parsers (Chen et al., 2023), and model counters (Usman et al., 2020). Research has shown the usefulness of differential testing in large-scale industrial software projects(Gulzar et al., 2019; Jarman et al., 2020a, b).
Silva et al. (Da Silva et al., 2020) study the feasibility of detecting semantic conflicts at merge time with differential testing. They use EvoSuite (Fraser and Arcuri, 2011) and Randoop (Pacheco and Ernst, 2007) to generate tests before a branch is merged. They show the generated tests can reveal semantic conflicts between the base and merging branches.
Recently, differential testing has also been used for testing deep learning programs and networks (Zhang et al., 2021). Guo et al. (Guo et al., 2018) propose Dlfuzz, a tool that generates rare inputs for DL networks. Dlfuzz uses a mutation algorithm to generate new inputs that activate previously unactivated neurons. Liu et al. (Liu et al., 2023a) propose a novel technique, called Gandalf, that generates inputs to detect inconsistencies between DL libraries.
@@
8. Conclusion
In this paper, we introduce Mokav, a novel tool that generates difference exposing tests (DET) employing LLMs with an execution-driven iterative approach. Mokav takes two versions of a program (P & Q) and an example test that produces the same output on P & Q. Mokav’s outcome is a test input that produces different results on P & Q. Mokav works with advanced prompting of an LLM: the prompt consists of P & Q, the example test, the output produced by the example test, and monitoring data collected during the execution of the example test. We evaluate Mokav on QuixBugs and a novel benchmark called C4DET, both in Python. Our experiments reveal that Mokav is effective and generates a DET for 81.7% (1,255/1,535) of program pairs in C4DET and 100% (32/32) of program pairs in QuixBugs. Mokav significantly outperforms Pynguin (Lukasczyk and Fraser, 2022) and DPP (Li et al., 2023), the closest, state-of-the-art test generation tools for the considered task.

### v2

@@ 6.1 Threats to Construct Validity
The main threat to the construct validity of this study is data leakage in LLMs [28]. As LLMs are trained on a large corpus of open-source data, they may have some program pairs of our datasets in their training data. This may threaten the generalizability of our experimental conclusions. Despite this concern, we note that open-source repositories usually do not put difference-exposing tests next to programs. This means there is no direct connection between programs and DETs in LLMs’ training data. Consequently, the data leakage problem is less significant in DET generation, the goal of Mokav, compared to more traditional tasks, such as code generation [29] and program repair [30]. Also, to further alleviate the data leakage concern, we assess the effectiveness of Mokav on two different datasets, i.e., QuixBugs [15] and C4DET.
6.2 Threats to Internal Validity
@@
Mokav uses a high temperature for the LLM in its default configuration. This leads to randomness in its responses, a threat to the internal validity of our experimental results. To address this threat, we repeat our main experiment of generating DETs for program pairs in C4DET five times. The results show that in all five instances, the effectiveness of Mokav is between 80.1% and 81.7%. This indicates that the main conclusions of our study are robust and only slightly affected by the randomness of LLMs.
6.3 Threats to External Validity
In this study, we evaluated Mokav on a dataset of program pairs with three constraints: (1) program length, (2) program language, and (3) program input types. These constraints might limit the generalizability of our results. Now we discuss each constraint, its implications, and potential solutions.
6.3.1 Program Length
As explained in Section 4.2, the program pairs in our dataset contain at most 5,000 tokens, 2,500 for each version. We consider this limit for our dataset to avoid exceeding the context size of widely-used LLMs, such as GPT-3.5 with a context size of 16,000 tokens. This decision might threaten the generalizability of our results to real-world programs. We believe it is not a blocker for two reasons. First, previous studies show that a large number of important code changes, such as bug fixes, occur at the single-function level [31], which are often small. This indicates the practical usefulness of generating DETs for program pairs with single-function diffs, which mostly have fewer than 5,000 tokens. Second, recent work highlights the possibility of using LLMs on coding tasks with a divide-and-conquer strategy [32]. This suggests that one could generate DETs for large programs by first splitting them into smaller units, trying to find a DET for one of the smaller units, and then assembling the results to obtain a DET for the full program. Using Mokav as a core component in a divide-and-conquer strategy for generating DETs for large programs is a promising direction for future work.
6.3.2 Program Language
In this study, we evaluate our approach on a dataset of Python programs. We focus on Python programs as the closest related work [17, 25] and our main baseline (DPP [17]) also evaluate their tools on Python programs. To have a quantitative and fair comparison, we have to consider the same language. This poses a threat to the generalizability of our results to other programming languages. We note that recent studies have shown that LLM-based tools are capable of multi-language test generation [33]. This indicates the potential of extending Mokav to generate DETs for other programming languages. We believe that the main challenge for such an extension is collecting execution differences for other languages. There are existing tools for execution differencing in other programming languages, such as Java [23], but we still need to adapt these tools and integrate them into the Mokav workflow. We consider this an opportunity for future work.
6.3.3 Program Input Types
The current version of Mokav is focused on generating inputs of primitive types (i.e., integers, floats, characters, and strings) and lists of such primitive data. This is a threat to the applicability of our approach to real-world programs, as many real-world projects have inputs with complex data types. Nevertheless, we believe that our proposed approach has significant real-world applications for two reasons. First, among the program pairs taken from the QuixBugs [15] dataset (see Section 4.2), four program pairs implement graph algorithms, such as the breadth first search. To generate DETs for these pairs, Mokav must produce inputs of the complex Node type. For this purpose, Mokav had to be modified to support generating Node inputs. One of the authors performed this modification with less than an hour of engineering effort. This shows that Mokav provides a flexible framework that can be quickly extended to generate DETs for real-world programs that use complex data types. Second, we note that recent studies demonstrate the strength of LLMs in generating tests with complex data types [34]. This is promising for Mokav to produce DETs for real-world projects with non-primitive inputs. We consider such an improvement as one of the main paths for future work.
6.4 Safety of LLM-generated Tests
As the adoption of LLMs for coding is growing, researchers have raised concerns about the safety of LLM-generated code [35]. They note that code snippets generated by LLMs can contain vulnerabilities, such as incorrect exception handling or well-known CWEs [36]. Such safety concerns also apply to tools like Mokav that employ LLMs for test generation and execute all outputs. For example, the execution of an LLM-generated code modifying file permissions could compromise the tester system. Since Mokav generated tests just call the program pairs with various inputs and lack a large test body, they are less prone to introducing such security problems. However, we still need to cautiously run LLM-generated tests to avoid potential issues, using safeguards such as rule-based validation of generated tests [37] or running tests in isolated sandbox environments [38], like container-based sandboxes [39].
7 Related Work
7.1 Test Generation with LLMs
@@
LLMs are being applied on many software engineering tasks [40], and test generation is one of the major targets [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 14, 10, 53, 54, 55, 56, 57]. A test consists of multiple parts: an input to a program under test (PUT), a test setup that moves the PUT into a specific state, and an oracle that asserts the PUT correctly produces the expected outputs. LLMs are used to generate all these parts of tests.
@@
Deng et al. [58] introduce TitanFuzz. This tool employs LLMs to generate human-like code that tests the API of deep learning (DL) libraries. TitanFuzz uses LLMs to generate and mutate input DL programs for fuzzing. Xia et al. [11] propose Fuzz4All, an LLM-based fuzzer that generates input test programs for compilers. Fuzz4All iteratively prompts the LLM to create new test inputs. The authors show that Fuzz4All effectively generates test inputs for six different programming languages. While TitanFuzz and Fuzz4All generate input programs for a single system, Mokav generates test inputs that differentiate two given programs.
@@
LLMs have also been used for oracle generation [59, 60, 61, 62]. Dinella et al. propose TOGA [12]. TOGA uses a fine-tuned version of CodeBERT [63] to generate assertion candidates based on the PUT and test prefix and rank these candidate assertions. Liu et al. [64] show limitations of TOGA: its unrealistic assumption of the availability of a correct focal method, its ignorance of important evaluation metrics, such as precision, and the lack of a straightforward baseline. Zhang et al. [65] generate tests in which the oracle expects an exception. In contrast to these works, Mokav focuses on input generation instead of oracle generation.
@@
Li et al. [17] introduce differential prompting, an LLM-based tool that generates fault-inducing tests. This tool takes a buggy program. It first uses the LLM to generate an intention description for the program. Then, it generates a fixed version of the program. Next, it employs the LLM to generate a list of test inputs. Finally, a test input that leads to different results on the buggy and fixed versions is labeled as fault-inducing. AID [25] is another tool that follows the same procedure, while it uses the LLM to generate a test generator, instead of test inputs. These two tools are different from Mokav as they generate fault-inducing tests for a given buggy program. In contrast, Mokav directly generates difference exposing tests for two given programs.
@@
Liu et al. [7] utilize LLMs to amplify existing tests in the HumanEval dataset [66]. They generate new tests that expose differences between the dataset programs and incorrect programs generated for the code documentation. Their test generation method starts from test seeds generated by LLMs and then improves them by type-aware mutation. This method is different from Mokav, which uses an execution-driven iterative approach to improve initial tests generated by the LLM.
7.2 Generating Tests for Python
@@
Generating tests for dynamically typed programming languages is challenging [67]. Auger [68] is one of the first tools to address this task for Python. Auger uses Python tracer to collect the values with which target functions are called and the values that target functions return. The collected data is then used to generate unit tests.
@@
The state-of-the-art tool for Python test generation is Pynguin [16, 69, 70], which uses evolutionary algorithms to generate unit tests with high coverage. It uses type annotations and static analysis to infer the type of function parameters. Lukasczyk et al. [71] conduct a study to improve Pynguin’s initial type inference strategy and find a more suitable evolutionary algorithm: DynaMOSA [27].
@@
Schoofs et al. [21] introduce AmPyfier, which amplifies existing Python test suites. AmPyfier runs existing tests on the PUT and collects values during execution. These collected values are then added to the existing test as new assertions. AmPyfier also amplifies the inputs to PUT iteratively. To make the amplification efficient, at each iteration, AmPyfier only keeps the assertion and input amplifications that increase the mutation or code coverage score.
@@
7.3 Differential Testing
Differential testing is a well-established technique for detecting anomalies and defects in programs [5, 72, 73, 74, 6, 75, 76, 77]. It is employed to identify issues in many types of programs, such as robotic vehicle programs [78], game agents [2], spreadsheet applications [79], certificate parsers [80], and model counters [81]. Research has shown the usefulness of differential testing in large-scale industrial software projects [1, 3, 82].
Silva et al. [8] study the feasibility of detecting semantic conflicts at merge time with differential testing. They use EvoSuite [83] and Randoop [84] to generate tests before a branch is merged. They show that the generated tests can reveal semantic conflicts between the base and merging branches.
Noller et al. [85] introduce HyDiff, a hybrid approach that generates inputs for two Java bytecode programs to expose differences, including functional differences. HyDiff uses AFL fuzzer [86] to generate new inputs by mutating existing ones, and also employs symbolic execution to generate inputs that cover various branches. In contrast to HyDiff, Mokav takes the source of two Python programs as input and leverages LLMs for test generation.
Recently, differential testing has also been used for testing deep learning programs and networks [87]. Guo et al. [88] propose Dlfuzz, a tool that generates rare inputs for DL networks. Dlfuzz uses a mutation algorithm to generate new inputs that activate previously unactivated neurons. Liu et al. [89] propose a novel technique, called Gandalf, that generates inputs to detect inconsistencies between DL libraries.
@@ 7.3 Differential Testing
8 Conclusion
In this paper, we introduce Mokav, a novel tool that generates difference exposing tests (DET) employing LLMs with an execution-driven iterative approach. Mokav takes two versions of a program (P & Q) and an example test that produces the same output on P & Q. Mokav’s outcome is a test input that produces different results on P & Q. Mokav works with advanced prompting of an LLM: the prompt consists of P & Q, the example test, the output produced by the example test, and monitoring data collected during the execution of the example test. We evaluate Mokav on QuixBugs and a novel benchmark called C4DET, both in Python. Our experiments reveal that Mokav is effective and generates a DET for 81.7% (1,255/1,535) of program pairs in C4DET and 100% (32/32) of program pairs in QuixBugs. Mokav significantly outperforms Pynguin [16] and DPP [17], the closest, state-of-the-art test generation tools for the considered task.